RHIC Startup:
Need 1 hour to adjust tune and coupling correction. Currently it is being tuned by hand. BBQ system is not tracking the tune. Did not agree with ARTUS. BBQ BTF thinks the tune is one thing, and BBQ LOCK thinks it is something else. Larry Hoff and Al Dellapena need to get together and agree on what the problem is, and the approach to a solution.
Need to go to slow factor 1. Don Bruno needs 3-4 ramps, then Carl Schultheiss needs to replace the connector to the Yellow Quad quench ground current that is reporting a spurios current. Then he will need 1-2 ramps to see that it is working well. This requires 2 hours.
Chipmunk failure has put in all beamstops all the way back to Tandem.
Both rings are now ramping together. Blue transmission 74%, Yellow 55%.
Yellow orbit correction on the ramp not working well. Todd should measure the lattice. Gamma_T setup needs work. Yellow frequency not synchronized with Blue. RF group can look at this behind ramps.
Yun wants to take an hour at store to check coupling angle.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email:sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |